Given events G and H: P(G) = 0.45; P(H) = 0.25; P(H AND G) = 0.19.
Part (a)
Find P(H OR G).
Part (b)
Find the probability of the complement of event (H AND G).
Part (c)
Find the probability of the complement of event (H OR G).

Answer by ikleyn(52864)

.

(a)  P(H or G) = P(H) + P(G) - P(H and G) = 0.25 + 0.45 - 0.19.    ANSWER


     You complete calculations.




(b)  P = 1 - P(H and G) = 1 - 0.19.    ANSWER


     You complete calculations.




(c)  It is the difference 1 and the probability you found/calculated in n a).


     Compute it on your own.

Solved, answered/instructed and completed.